An object 3.3 feet tall casts a shadow that is 9.9 feet long. How long in feet would the shadow be for an object which is 18.3 feet tall?
Answer by TimothyLamb(4379)

---
3.3/9.9 = 18.3/s
s = 18.3/(3.3/9.9)
s = 54.9 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php